# December 2023 Social and Economic Impact Analysis of Solar Mini-Grids in Rural Africa: A Cohort Study from Kenya and Nigeria

### v1

@@ 1 Introduction
Sub-Saharan Africa (SSA) has vast renewable energy resources, including solar, wind, hydropower, and geothermal [1]. However, these resources remain largely untapped, and the region is still heavily reliant on fossil fuels due to factors including a lack of infrastructure, financing, and technical expertise [2]. Decentralized renewable energy (DRE) solutions can offer a promising alternative to grid-based electricity in SSA [3, 4]. DRE systems like the one illustrated in Fig. 1 are typically smaller in scale and can be installed in rural areas where grid extension is not feasible or cost-effective. DRE systems can also be operated by local communities, which can help to empower these communities and promote economic development [5, 6]. Several studies have investigated the potential of renewable mini-grids (RMGs) as decentralized renewable energy solutions in SSA [7, 8, 9]. These studies have found that RMGs can have a significant positive impact on access to electricity, livelihoods, and economic development [10]. RMGs are small power grids that typically serve a few hundred to a few thousand households and businesses. They are often powered by renewable energy sources, such as solar or wind, and can provide a reliable and affordable source of electricity to communities that are not connected to the national grid.
@@
Rural electrification can significantly benefit communities [17]. According to [16], the introduction of electrification via RMG has had significant positive impacts on socioeconomic factors such as health care and economic development. RMG systems promote faster and more flexible energy supply, especially when integrated with additional components such as electricity storage in a hybrid system [18]. The slow extension of conventional grid systems has led to increased interest in RMGs, a decentralized alternative not connected to the public grid and generating electricity based on various technologies [19]. The main obstacle that prevented private households from using electricity was the high price. However, according to [15], the incremental cost of providing electricity access to households is below the cost of alternatives such as kerosene lighting.
@@
While existing research has acknowledged the significance of social impacts stemming from RMG implementation in SSA, there remains a notable absence in empirical measurements [24]. The prevailing studies predominantly focus on technical and economic aspects, neglecting the quantification of social ramifications [25, 26, 27, 28]. A recent study in the Gbamu Gbamu village in Nigeria uses a combination of statistical tests to measure the financial impact of RMGs on local businesses using predominantly economic factors including gender, marital status, household size, age, education level, years of business establishment, hours of operation, building tenure, capital source, number of employees, generator ownership, and the days of operation [29].
@@
Our study aims to bridge the gap in empirical measurement of social impact of RMGs on rural communities by introducing a comprehensive framework that delineates five crucial axes encompassing gender equality, productivity, health, safety, and economic activity. By leveraging a survey designed specifically around these axes, we endeavor to quantify the multifaceted impacts of solar RMGs on various social and economic factors within rural African communities. Unlike previous single-region analyses, our research embraces a comparative approach, drawing insights from two distinct countries, Kenya and Nigeria. This deliberate expansion of our dataset promises a richer understanding of the nuanced effects of RMG interventions across diverse sociocultural landscapes. The rest of the paper is organized as follows. In Section 2, we discuss the model, data, and estimation techniques. The results and findings are presented in Section 3. Further discussion and methodology limitations are shown in Section 4. We conclude in Section 5.
@@ 2.1 Project Description and Data Collection
Between 2021 and 2023, we installed solar mini-grids in and conducted a study across 22 communities in Nigeria and Kenya to assess if there were any significant changes regarding the quality of life of the populations connected to their mini-grids. To collect the necessary information, we used surveys composed of semi-structured questionnaires that captured quantitative and qualitative data such as demographics, access to education for children, access to clean drinking water, creation of jobs and businesses, and economic opportunities for women, as well as many others. Primary data was collected from a variety of respondents categorized either as households or commercial and institutional organizations (e.g. businesses, schools, clinics, etc.) through face-to-face interviews by independent, third-party field enumerators.
This study examined the impact of the mini-grids on key performance indicators (KPIs) by comparing responses before and one year after installing a solar mini-grid in the respective communities. Due to the lack of data from communities without a solar mini-grid, only pre-and post-treatment data have been used in this study. The pre-treatment survey was conducted between 2021 and 2022, while the post-treatment survey was conducted between 2022 and 2023.
@@ 2.2 Survey Design
Surveys were carried out in person by independent, third-party surveyors who used electronic forms to minimize transcription bias. Participants were briefed and consent obtained through signed forms, ensuring ethical compliance. No incentives were offered.
@@ 2.3.1 Paired Samples t-Test
where is the mean value of the differences between paired samples, the standard deviation the differences between paired samples, and is the sample size.
@@ 2.3.2 Wilcoxon Signed-Rank Test
The Wilcoxon signed-rank test is a statistical method to compare two dependent samples from matched or paired data. While it does assume the distribution of the differences is symmetric, it does not assume any specific distribution of the samples themselves and serves as a non-parametric equivalent to the paired samples -test, particularly applicable to categorical variables with meaningful differences between ranks (i.e. ordinal data). The test is evaluated taking into account both the sign and magnitudes of observed differences. The Wilcoxon signed-rank test is implemented using the null and alternative hypotheses:
@@ 2.3.6 Linear Regression t-Test
where is the expected value of response; is the intercept, i.e the expected value of response when the predictor is 0; is the slope coefficient, i.e the average change in the response given a unit increase in the predictor; and is the value of predictor.
@@ 3.1 Survey Distribution
There are a total of 86 questions for the pre-connection household survey and 93 and 37 for the post-connection household and organizational surveys, respectively. The pre- and post-connection household surveys shared 64 questions, thus enabling direct comparison of individual customer responses before and after connection. The questions were distributed among the five KPIs according to Fig. 2. Correlation tables detailing the responses to the survey questions are presented in Figs. 14, LABEL:, 15, LABEL:, and 16 in the appendix.
@@ 3.3 Gender Equality
The full results of the statistical analysis for questions pertaining to gender equality are available in Table 4 in the appendix.
@@ 3.4 Productivity
The full results of the statistical analysis for questions pertaining to productivity are available in Table 5 in the appendix.
@@
Prior to connecting to the mini-grid, a majority of the households had limited or no reliable power sources. Specifically, 55% of households depended on petrol generators. However, one year after the mini-grid connection, this number dramatically decreased to just 1%, with the mini-grid becoming the primary power source for 91% of survey respondents. The shift was also evident in the broader reduction of unclean energy sources, including diesel, petrol, and kerosene, which decreased from 66% in the pre-survey to 19% in the post-survey. Among paired households, this represented a 7 percentage-point reduction. As illustrated in Fig. 6(a), every 100 additional customers connected onto the mini-grid would result in fewer households using unclean power sources.
@@ 3.5 Health
The full results of the statistical analysis for questions pertaining to health are available in Table 6 in the appendix.
@@
The transition to the mini-grid has led to notable improvements in water quality for the surveyed communities. There was a decrease in the use of dirty water, with pre-survey respondents reporting a 38% usage rate, which dropped to just 15% in the post-survey. Concurrently, the majority of households shifted to using community wells or pumps, accounting for 50% of the water sources post-connection.
@@ 3.6 Safety
The full results of the statistical analysis for questions pertaining to safety are available in Table 7 in the appendix.
@@ 3.7 Economic Activity
The full results of the statistical analysis for questions pertaining to economic activity are available in Table 8 in the appendix.
@@ 4 Discussion and Qualitative Insights
Before the mini-grid connection, many economically productive activities were often determined by the natural light hours and were limited by little-to-no access to electricity. For instance, many residents had to travel to charge electronic appliances such as phones and, so, restrict the time they could spend on other activities. Respondents involved in the fishing industry noted that their fishing activities were limited due to the absence of ice for preserving their catch before reaching shore. Consequently, they restricted their daily catch to quantities they could sell immediately, as smoking was the only available method for preservation. These inefficiencies were curtailed or eliminated by a reliable power source. More broadly, productivity is related to light hours for households, time spent on and distance traveled for specific activities like water collection and appliance charging, and enhanced operational hours and workforce capacity for organizations. The surveys contained 31 productivity-related questions for households and 7 for businesses. Analyses of the metrics related to productivity demonstrated gains and savings in terms of resources allocated to various activities, thus pointing to greater efficiency at the individual and community levels. Connecting to the mini-grid gave more light hours, reduced reliance on unclean or non-renewable power sources, and provided on-demand charging stations at home for all their electronic devices. The responses from open-ended survey questions provided additional depth and nuance to these findings. The availability of electronic information devices such as televisions and radios has expanded access to information for residents, while the convenience of charging phones has enhanced communication within families and the community at large. The ability to access social media platforms has also heightened residents’ awareness of local, national, and international news. Furthermore, the acquisition of electric appliances like clothes irons has enabled residents to wear ironed clothes, contributing to an improved sense of self-presentation. There was also an improvement in the academic performance of school-aged children affirmed by both households and schools. Children could now extend their study hours at home, and schools could provide more services to support the student body, potentially at later hours. Most organizations also had a notable increase in their hours of operation—now with a reliable source of electricity—thus driving more business growth in hiring more workers and expanding their lines of products and services.
@@
Sixteen health-related questions were posed to households, focusing on the ease of access to the nearest clinic, use of kerosene lamps, water purification methods, and the overall quality of healthcare. The results highlighted a significant positive impact on the health and well-being of communities connected to mini-grids. Residents gained access to clean drinking water, primarily through community wells or pumps, reducing health risks associated with the use of unclean water. Clinics reported shorter wait times, extended hours of operation, the capacity to treat more patients, and improved cold storage for vaccines and medicines. Residents directly confirmed an overall enhancement in healthcare services and living conditions. The adoption of mini-grids led many households to discontinue the use of hazardous kerosene lamps, known for their risks of poisoning, fires, and explosions. Here, too, open-ended survey responses further enriched the findings. The consistent voltage supply from the mini-grid has enabled a shift away from petrol-powered generators, significantly reducing noise and pollution in residential areas. Residents, like clinics, also benefit from cold storage facilities, allowing access to cold drinking water and other chilled beverages as needed, alongside a notable reduction in food wastage. This transition has contributed to an improved quality of life for the residents, underscoring the multifaceted health and lifestyle benefits brought about by the mini-grid installation.
@@
Survey questions measuring economic activity assessed changes in household incomes and spending, earnings for businesses, employment, and business ownership. This section contained 17 questions. The median household income in Nigeria decreased between 2021 and 2023, but this can likely be best explained by macroeconomic trends [34, 35]. On the other hand, Kenyan households observed an increase in income after mini-grid connection. Among the paired respondents—those community members who were present both before the mini-grid installation and at least until the post-survey one year after connection—there was a significant rise in median monthly income. This group saw their median income increase from 4,000 KES to 18,000 KES, highlighting a substantial economic uplift due to the mini-grid installation. Across both countries, business ownership was higher, and previously established businesses had increased earnings. The mini-grid installation bolstered the entrepreneurial economy as electrification brought additional economic opportunities. Households could also save on the charging cost for phones and appliances, cooking energy, and water costs, freeing up money for other investments.
@@ 5 Conclusion
This paper offers an in-depth evaluation of the tangible socioeconomic impacts of solar mini-grids in rural communities across sub-Saharan Africa. It utilizes robust empirical methods to examine five Key Performance Indicators (KPIs): gender equality, productivity, health, safety, and economic activity. This constitutes the first comprehensive analysis of the social and economic effects of solar mini-grids in rural African settings.
@@
Despite the study’s positive findings, certain limitations were encountered. The survey’s structure led to an imbalance between discrete and continuous variables, limiting the pool available for normal distribution in regression analyses. This restriction affected the robustness of testing electricity consumption effects at the household level. Site-level analyses provided some compensation, but they offered only community-wide results, amalgamating individual household and business impacts. Furthermore, many survey questions yielded outputs unsuitable for parametric hypothesis testing due to their lack of meaningful order or magnitude.
@@
Solar mini-grids have significantly driven positive transformations in communities, leading to enhanced opportunities for women and girls, better healthcare, and economic growth. Their impact on various socioeconomic factors highlights their role in achieving the Sustainable Development Goals (SDGs) by 2030.
@@
This project was primarily funded by Renewvia Energy Corporation. While no specific grants were allocated for this project, authors employed by Renewvia Energy Corporation led the study design, analysis, interpretation, and writing of the report. This collaboration was in partnership with the African Leadership University, the University of Nairobi, and UC Berkeley. Data collection was carried out by independent, third-party surveyors.
@@
In line with our commitment to transparency and scientific reproducibility, the anonymized responses from the surveys conducted for this study are released alongside the publication of this paper. We have ensured that all personally identifiable information has been removed from these responses to maintain the confidentiality and privacy of the participants. This anonymization process was undertaken with great care to preserve the integrity and utility of the data for further analysis and research.

### v2

@@ 1 Introduction
Sub-Saharan Africa (SSA) has vast renewable energy resources, including solar, wind, hydropower, and geothermal [1]. However, these resources remain largely untapped, and the region is still heavily reliant on fossil fuels due to factors including a lack of infrastructure, financing, and technical expertise [2]. Decentralized renewable energy solutions, especially renewable mini-grid (RMG) systems, can offer a promising alternative to grid-based electricity in SSA [3, 4]. RMG systems like the one illustrated in Fig. 1 are typically smaller in scale and can be installed in rural areas where grid extension is not feasible or cost-effective. RMG systems can also be operated by local communities, empowering these communities and promoting economic development [5, 6]. Several studies have investigated the potential of RMGs as decentralized renewable energy solutions in SSA [7, 8, 9]. These studies have found that RMGs can have a significant positive impact on access to electricity, livelihoods, and economic development [10]. RMGs are small power grids that typically serve a few hundred to a few thousand households and businesses. Often powered by renewable energy sources, such as solar or wind, they can provide a reliable and affordable source of electricity to communities that are not connected to the national grid.
@@
Rural electrification can significantly benefit communities [17]. According to [16], the introduction of electrification via RMG has had significant positive impacts on socioeconomic factors such as health care and economic development. RMG systems promote faster and more flexible energy supply, especially when integrated with additional components such as electricity storage in a hybrid system [18]. The slow extension of conventional grid systems has led to increased interest in RMGs, a decentralized alternative not connected to the public grid and generating electricity based on various technologies [19]. According to [15], the incremental cost of providing electricity access to households is below the cost of alternatives such as kerosene lighting.
@@
While existing research has acknowledged the significance of social impacts stemming from RMG implementation in SSA, there remains a notable absence in empirical measurements [24]. The prevailing studies predominantly focus on technical and economic aspects, neglecting the quantification of social ramifications [25, 26, 27, 28]. A recent study in the Gbamu Gbamu village in Nigeria uses statistical tests to measure the financial impact of RMGs on local businesses using predominantly economic factors including gender, marital status, household size, age, education level, years of business establishment, hours of operation, building tenure, capital source, number of employees, generator ownership, and the days of operation [29].
@@
Our study aims to bridge the gap in empirical measurement of social impact of RMGs on rural communities by introducing a comprehensive framework that delineates five crucial axes encompassing gender equality, productivity, health, safety, and economic activity. By leveraging a survey designed specifically around these axes, we endeavor to quantify the multifaceted impacts of solar RMGs on various social and economic factors within rural African communities. Unlike previous single-region analyses, our research embraces a comparative approach, drawing insights from two distinct countries, Kenya and Nigeria. This deliberate expansion of our dataset yields a richer understanding of the nuanced effects of RMG interventions across diverse sociocultural landscapes. The rest of the paper is organized as follows. In Section 2, we discuss the model, data, and estimation techniques. The results and findings are presented in Section 3. Further discussion and methodology limitations are shown in Section 4. We conclude in Section 5.
@@ 2.1 Project Description and Data Collection
Between 2021 and 2023, we installed solar mini-grids in and conducted a study across 22 communities in Nigeria and Kenya to assess if there were any significant changes regarding the quality of life of the populations connected to their mini-grids. The capacity of the mini-grids varied, starting from 6.24 kWp with 14.8 kWh of battery storage, up to 541 kWp accompanied by 1.10 MWh of battery storage. To collect the necessary information, we used surveys composed of semi-structured questionnaires that captured quantitative and qualitative data such as demographics, access to education for children, access to clean drinking water, creation of jobs and businesses, and economic opportunities for women, as well as many others. Primary data was collected from a variety of respondents categorized either as households or commercial and institutional organizations (e.g. businesses, schools, clinics, etc.) through face-to-face interviews by independent, third-party field enumerators.
This study examined the impact of the mini-grids on the key performance indicators (KPIs) by comparing responses before and one year after installing a solar mini-grid in the respective communities. Due to the lack of data from communities without a solar mini-grid, only pre-and post-treatment data have been used in this study. The pre-treatment survey was conducted between 2021 and 2022, while the post-treatment survey was conducted between 2022 and 2023.
@@ 2.2 Survey Design
Surveys were carried out in person by independent, third-party surveyors who used CommCare electronic forms to minimize transcription bias. Participants were briefed and consent obtained through signed forms, ensuring ethical compliance. No incentives were offered.
@@ 2.3.1 Paired Samples t-Test
where is the mean value of the differences between paired samples, the standard deviation of the differences between paired samples, and is the sample size.
@@ 2.3.2 Wilcoxon Signed-Rank Test
The Wilcoxon signed-rank test is a statistical method to compare two dependent samples from paired data. While it assumes the distribution of the differences is symmetric, it does not assume any specific distribution of the samples themselves and serves as a non-parametric equivalent to the paired samples -test, particularly applicable to categorical variables with meaningful differences between ranks (i.e. ordinal data). The test is evaluated taking into account both the sign and magnitudes of observed differences. The Wilcoxon signed-rank test is implemented using the null and alternative hypotheses:
@@ 2.3.6 Linear Regression t-Test
where is the expected value of response; is the intercept, i.e. the expected value of response when the predictor is 0; is the slope coefficient, i.e. the average change in the response given a unit increase in the predictor; and is the value of predictor.
@@ 3.1 Survey Distribution
There are a total of 86 questions for the pre-connection household survey and 93 and 37 for the post-connection household and organizational surveys, respectively. The pre- and post-connection household surveys shared 64 questions, enabling direct comparison of individual customer responses before and after connection. The questions were distributed among the five KPIs according to Fig. 2. Correlation tables of the responses are presented in Figs. 14, LABEL:, 15, LABEL:, and 16 in the appendix.
@@ 3.3 Gender Equality
The full results of the statistical analysis for questions pertaining to gender equality are available in LABEL:tab:app:gender in the appendix.
@@ 3.4 Productivity
The full results of the statistical analysis for questions pertaining to productivity are available in LABEL:tab:app:productivity in the appendix.
@@
Prior to connecting to the mini-grid, a majority of the households had limited or no reliable power sources. Specifically, 55% of households depended on petrol generators. However, one year after the mini-grid connection, this number dramatically decreased to just 1%, with the mini-grid becoming the primary power source for 91% of survey respondents. The shift was also evident in the broader reduction of unclean energy sources, including diesel, petrol, and kerosene, which decreased from 66% in the pre-survey to 19% in the post-survey. Among paired households, this represented a 7 percentage-point reduction. As illustrated in Fig. 7(a), every 100 additional customers connected to the mini-grid would result in fewer households using unclean power sources.
@@ 3.5 Health
The full results of the statistical analysis for questions pertaining to health are available in LABEL:tab:app:health in the appendix.
@@
The installation of mini-grids has led to notable improvements in water quality for the surveyed communities. There was a decrease in the use of dirty water, with pre-survey respondents reporting a 38% usage rate, which dropped to just 15% in the post-survey. Concurrently, the majority of households shifted to using community wells or pumps, accounting for 50% of the water sources post-connection.
@@ 3.6 Safety
The full results of the statistical analysis for questions pertaining to safety are available in LABEL:tab:app:safety in the appendix.
@@ 3.7 Economic Activity
The full results of the statistical analysis for questions pertaining to economic activity are available in LABEL:tab:app:econ in the appendix.
@@ 4 Discussion and Qualitative Insights
Before the mini-grid connection, many economically productive activities were often determined by the natural light hours and were limited by little-to-no access to electricity. For instance, many residents had to travel to charge electronic appliances such as phones and, so, restrict the time they could spend on other activities. Respondents involved in the fishing industry noted that their fishing activities were limited due to the absence of ice for preserving their catch before reaching shore. Consequently, they restricted their daily catch to quantities they could sell immediately, as smoking was the only available method for preservation. These inefficiencies were curtailed or eliminated by a reliable power source. More broadly, productivity is related to light hours for households, time spent on and distance traveled for specific activities like water collection and appliance charging, and enhanced operational hours and workforce capacity for organizations. The surveys contained 31 productivity-related questions for households and 7 for businesses. Analyses of the metrics related to productivity demonstrated gains and savings in terms of resources allocated to various activities, thus pointing to greater efficiency at the individual and community levels. Connecting to the mini-grid gave more light hours, reduced reliance on unclean or non-renewable power sources, and provided on-demand charging stations at home for all their electronic devices. This result affirms the conclusions of [26, 27], where the installations of solar panels and a small hydroelectric power plant in Ethiopian and Rwandan communities, respectively, led to increased employment, educational attainment, and business activity. Our study expands upon this result by examining a broader set of questions that capture economic productivity both at the household and business level. The responses from open-ended survey questions provided additional depth and nuance to these findings. The availability of electronic information devices such as televisions and radios has expanded access to information for residents, while the convenience of charging phones has enhanced communication within families and the community at large. The ability to access social media platforms has also heightened residents’ awareness of local, national, and international news. Furthermore, the acquisition of electric appliances like clothes irons has enabled residents to wear ironed clothes, contributing to an improved sense of self-presentation. There was also an improvement in the academic performance of school-aged children affirmed by both households and schools. Children could now extend their study hours at home, and schools could provide more services to support the student body, potentially at later hours. Most organizations also had a notable increase in their hours of operation—now with a reliable source of electricity—thus driving more business growth in hiring more workers and expanding their lines of products and services.
@@
Sixteen health-related questions were posed to households, focusing on the ease of access to the nearest clinic, use of kerosene lamps, water purification methods, and the overall quality of healthcare. The results highlighted a significant positive impact on the health and well-being of communities connected to mini-grids. Residents gained access to clean drinking water, primarily through community wells or pumps, reducing health risks associated with the use of unclean water. Clinics reported shorter wait times, extended hours of operation, the capacity to treat more patients, and improved cold storage for vaccines and medicines. Residents directly confirmed an overall enhancement in healthcare services and living conditions. The adoption of mini-grids led many households to discontinue the use of hazardous kerosene lamps, known for their risks of poisoning, fires, and explosions. Here, too, open-ended survey responses further enriched the findings. The consistent voltage supply from the mini-grid has enabled a shift away from petrol-powered generators, significantly reducing noise and pollution in residential areas. Residents, like clinics, also benefit from cold storage facilities, allowing access to cold drinking water and other chilled beverages, alongside a notable reduction in food wastage. This transition has contributed to an improved quality of life for the residents, underscoring the multifaceted health and lifestyle benefits brought about by the mini-grid installation.
@@
Survey questions measuring economic activity assessed changes in household incomes and spending, earnings for businesses, employment, and business ownership. This section contained 17 questions. The median household income in Nigeria decreased between 2021 and 2023, but this can likely be best explained by macroeconomic trends [34, 35]. On the other hand, Kenyan households observed an increase in income after mini-grid connection. Among the paired respondents—those community members who were present both before the mini-grid installation and at least until the post-survey one year after connection—there was a significant rise in median monthly income. This group saw their median income increase from 4,000 KES to 18,000 KES, highlighting a substantial economic uplift due to the mini-grid installation. This is in contrast to the findings of [28], which did not find meaningful impacts on economic and non-economic outcomes after the expansion of electric grid infrastructure in rural Kenya, potentially due to the low energy utilization they observed at the household level.
Across both countries, business ownership was higher, and previously established businesses had increased earnings. The mini-grid installation bolstered the entrepreneurial economy as electrification brought additional economic opportunities. Households could also save on the charging cost for phones and appliances, cooking energy, and water costs, freeing up money for other investments.
@@ 5 Conclusion
This paper offers an in-depth evaluation of the tangible socioeconomic impacts of solar mini-grids in rural communities across sub-Saharan Africa. It utilizes robust empirical methods to examine five Key Performance Indicators: gender equality, productivity, health, safety, and economic activity. This constitutes the first comprehensive analysis of the social and economic effects of solar mini-grids in rural African settings.
@@
Despite the study’s positive findings, certain limitations were encountered. The survey’s structure led to an imbalance between discrete and continuous variables, limiting the use of regression analyses. This restriction affected the robustness of testing electricity consumption effects at the household level. Site-level analyses provided some compensation, but they offered only community-wide results, amalgamating individual household and business impacts. Furthermore, many survey questions yielded outputs unsuitable for parametric hypothesis testing due to their lack of meaningful order or magnitude.
@@
Solar mini-grids have significantly driven positive transformations in communities, leading to enhanced opportunities for women and girls, better healthcare, and economic growth. Their impact on various socioeconomic factors highlights their role in achieving the Sustainable Development Goals by 2030.
@@
This project was primarily funded by Renewvia Energy Corporation. While no specific grants were allocated for this project, authors employed by Renewvia Energy Corporation led the study design, analysis, interpretation, and writing of the report. This collaboration was in partnership with the African Leadership University; the University of Nairobi; and the University of California, Berkeley. Data collection was carried out by independent, third-party surveyors.
@@
In line with our commitment to transparency and scientific reproducibility, the anonymized responses from the surveys conducted for this study are released alongside the publication of this paper. We have ensured that all personally identifiable information has been removed from these responses to maintain the confidentiality and privacy of the participants. Per the request of the communities, this includes redacting the names and precise locations of the communities involved, replacing them with the state in which each community resides (e.g, “Cross River 1,” “Turkana 2,” etc.). This measure is crucial to protect the identity of the communities while still providing context and geographical relevance to the data. This anonymization process was undertaken with great care to preserve the integrity and utility of the data for further analysis and research.